OpenStudy (anonymous):

Write an equation of the line that passes through (-5,-1) and is parallel to the line .y= 4x-6

hartnn (hartnn):

@Chikibabylovesyou Hi, \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) can you find the slope of line y=4x-6 ?

OpenStudy (anonymous):

The slope is 4

hartnn (hartnn):

parallel lines have same slope, so the slope of required line is also 4. now you have slope = m=4 and a point = (x1,y1)=(-5,-1) can you find the equation of line ?

hartnn (hartnn):

the point slope form of equation of line is \(\large y-y_1=m(x-x_1)\) where m=slope =4 here. point =(x1,y1)= (-5,-1) --->x1=-5,y1=-1 just plug in values!

OpenStudy (anonymous):

So, y=4x+19

hartnn (hartnn):

yes, thats correct! good work :)

OpenStudy (anonymous):

Thank you so much (: